The sum of two integers is eighteen. Eight times the smaller integer is twenty-four more than two times the larger integer. Find the integers.
Answer by oberobic(2304)

x = lesser integer
y = greater integer
.
x + y = 18
8x = 2y +24
.
You can solve by substitution or by systems of linear equations.
.
x + y = 18
8x - 2y = 24
.
2x + 2y = 36
8x - 2y = 24
10x = 60
x = 6
.
6+y = 18
y = 12
.
Check your work!
.
8x = 8*6 = 48
2y = 2*12 = 24
Correct.
.
Answer:
The integers are 6 and 12.
.
Done.